OpenStudy (anonymous):

1. The sum of two rational numbers is 11/16, and one of them is 5/6 of the other. Find the rational numbers. 2. The numerator of a rational number exceeds it denominator by 3. If 1 is added to the denominator and 5 is subtracted from the numerator, the value of the rational number becomes 1/2. What is the rational number?

OpenStudy (shubhamsrg):

1) seems tough!1 hmm..

OpenStudy (anonymous):

1)answer is -7/48

hero (hero):

x + 5/6 = 11/16

hero (hero):

Give help, not answers....that's like rule number one of Open Study

OpenStudy (anonymous):

how do you solve it?

OpenStudy (anonymous):

x = 66/80 buy cross multiply x = 33/40

OpenStudy (anonymous):

@Hero what about another one

Directrix (directrix):

1. x + 5/6 = 11/16 x = 11/16 - 5/6 x = 33/48 - 40/48 x = -7/48 -->